OpenStudy (anonymous):

critical point of f(x) -3/x^2+9

OpenStudy (anonymous):

Derivative set equal to zero and solve.

OpenStudy (anonymous):

The derivative is\[-3 (x ^{2}+9)(2x)\]

OpenStudy (anonymous):

x^2+9 is the denominator?

OpenStudy (anonymous):

Yes, to the -2, forgot to add that.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use the quotient rule

OpenStudy (anonymous):

6x/(x^2+9x)2??

OpenStudy (anonymous):

i mean 6x/(x^2+9x)^2?

OpenStudy (anonymous):

\[f'(x)=\frac{-3(2x)}{(x^2+9)^2}=\frac{-6x}{(x^2+9)^2}\]

OpenStudy (anonymous):

@Dockworker has it

OpenStudy (anonymous):

actually, numerator is positive

OpenStudy (anonymous):

\[f"(x)=\frac{6x}{(x^2+9)^2}\]

OpenStudy (anonymous):

Yes, forgot the negative in front of the 3

OpenStudy (anonymous):

f'(x), not f''(x) rather :)

OpenStudy (anonymous):

set that equal to zero?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes, normally for critical numbers you'd find the zeros and when the derivative is undefined, but because the denominator can never be 0 in this case, there's only one critical number

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

If you plug it into Wolfram Alpha you can check your answer.

OpenStudy (anonymous):

they gave me 0

OpenStudy (anonymous):

thank you